OpenStudy (druidsoul):

Determine the projection of u=1.6i + 3.3j in the v=-2.1i -0.5j direction.

OpenStudy (anonymous):

\[ projection~of ~u~on~v=\frac{ u.v }{ \left| v \right| }\] \[vector~projection ~of~u~on~v~\in~the~direction~of ~v=\frac{ u.v }{ \left| v \right| }(\frac{ v }{ \left| v \right| })\]

OpenStudy (mathmale):

Yes, there are both vector and scalar projections. Their major difference is that one is a scalar quantity (no direction) and the other is a vector quantity (has direction indicated by the unit vector \[\frac{ v }{ |v| }\]